OpenStudy (anonymous):

Factor completely: -3x^2 + 6x - 9

OpenStudy (anonymous):

= -3(x^2 + 2x - 3) = -3(x-1)(x+3)

OpenStudy (anonymous):

@scarydoor if you removed -3 as the GCF, how come you left -3? Can this be factored further?

OpenStudy (anonymous):

oh whoops, I did the thing above wrong: -3x^2 + 6x -9 = -3(x^2 -2x + 3) = ... something.

OpenStudy (anonymous):

Can't be factored further unless you want to use complex numbers.